Countryyy:

Two events, A and B, are associated with the following probabilities: P(A)=2/3 P(A∩B)=2/9 If the events are independent, what is the probability of B? P(B)= __[blank]__ Enter your answer as the fraction that correctly fills in the blank, using the / symbol, like this: 42/53

Zarkon:

If A and B are independent then \(P(A\cap B)=P(A)P(B)\) then \[\frac{2}{9}=\frac{2}{3}P(B)\] solve for \(P(B)\)

joshiscool:

Zarkon Where art thou Zarkon

Allison:

If you're question is answered, please close it. Thank you.